OpenStudy (anonymous):

How to approximate a number using f(x)=-1/4x+3/2

OpenStudy (anonymous):

It wants the approx of \[\sqrt{3.9}\] and i know i need to plug it in for x but i dont know how

OpenStudy (anonymous):

It wants the approximate of it, so i cant just plug in \[\sqrt{3.9}\]. i should get the decimal approx of it using this formula.

OpenStudy (anonymous):

A linear approximation... like this?\[ f(x_2)=f'(x_1)(x_2-x_1)+x_1 \]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Let \(x_1=4\) since it's easy to find the square root of it, and it's really close to \(3.9\)

OpenStudy (anonymous):

Should have written: \[ f(x_2)\approx f'(x_1)(x_2-x_1)+f(x_1) \]

OpenStudy (anonymous):

Can i plug it in for x using the simplified form?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

First of all, what is \(f'(x)\) in this case?

OpenStudy (anonymous):

is there any reason to put 2-(1/4)(x+2) into f(x)=-1/4x+3/2

OpenStudy (anonymous):

I have no idea wtf you're doing. You come outta nowhere with these equations.

OpenStudy (anonymous):

f'(a)= -(1/4) f(a)= 2

OpenStudy (anonymous):

sorry lol. Im doing an linear approx of f(x)= radical(2-x) at a=-2

OpenStudy (anonymous):

Okay, you just need to do \(x_2-x_1\)... What is the problem?

OpenStudy (anonymous):

I have it written as f(x)=f(a)+f'(a)(x-a)

OpenStudy (anonymous):

So you're using a linear approximation at \(4\). That is a statement. What is the question?

OpenStudy (anonymous):

I dont understand what you are saying.

OpenStudy (anonymous):

I really am trying. its just not making sense. you are plugging in 4 for f(a)?

OpenStudy (anonymous):

no that cant be right. it should be plugged in for x

OpenStudy (anonymous):

I just cant get the right approx when i plug 4 in for x.

OpenStudy (anonymous):

maybe i did the values wrong.

OpenStudy (anonymous):

\[ f(x_2)\approx f'(x_1)(x_2-x_1)+f(x_1) \]We have \(x_1=4,x_2=3.9\) \[ f(3.9)\approx f'(4)(3.9-4)+f(4) \]

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

\[f(3.9)=3.6\]

OpenStudy (anonymous):

its i got it. you just take 4-3.9= 0.1 then you plug it in.